OpenStudy (anonymous):

Find a quadratic equation equation whose roots differ by 1 and are reciprocals of each other. Find the roots of the equation.

OpenStudy (anonymous):

So you're solving the system a-b=1 ab=1 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

What do you have so far? That system yields a quadratic itself.

OpenStudy (anonymous):

If a-b=1, then a=b+1 [a is larger root, b is smaller root) -> ab=1=(b+1)b -> b^2+b-1=0. Can use quadratic formula to solve for b, then add 1 to that to get a. With the two roots known, you can find a polynomial.